### v1

# The equivalence of asymptotic average shadowing and vague specification properties and its consequences

Abstract
We establish the equivalence between the asymptotic average shadowing and the vague specification properties and we use this equivalence to answer a question posed by Downarowicz and Weiss [Ergod. Th. & Dynam. Sys. 44(9) (2024), 2565-2580]. Additionally, we prove that the asymptotic average shadowing property is equivalent to the average shadowing property if the phase space is complete with respect to the dynamical Besicovitch pseudometric. We use the latter result to prove that proximal and minimal shift spaces from [Minimal and proximal examples of $\bar{d}$-stable and $\bar{d}$-approachable shift spaces, Ergod. Th. & Dynam. Sys., (2024) First View] possess the asymptotic average shadowing property.
1 Introduction
00footnotetext: Mathematics Subject Classification 2020. Primary 37B65 ; Secondary 37B05, 37B1000footnotetext: Key words. shadowing property, specification property, Besicovitch pseudometric, shift space.
The shadowing property, also called the pseudo-orbit tracing property, is a fundamental concept in the theory of dynamical systems. Informally, it means that every pseudo-orbit (a sequence of points in the phase space such that every point is close to the image of the previous one) is traced by an actual orbit. Such property played a crucial role in the development of the theory of hyperbolic dynamical systems because of the Shadowing Lemma which is a key step in the proof of structural stability of hyperbolic sets (see [4, 12]).
Besides its applications to differentiable dynamics, the shadowing property became a subject of studies in topological dynamics. This led to an introduction of different notions of pseudo-orbits and tracing, resulting in variants of the shadowing property, such as the limit shadowing property [8], the average shadowing property [3], and the asymptotic average shadowing property [9].
In the early 70’s Bowen [4] introduced another notion of tracing of sequence of points in the space by an orbit. Bowen called it the specification property, and used it to study Axiom $A$ diffeomorphisms. Roughly speaking, the specification property guarantees the existence of an orbit to trace arbitrary, but finite, collection of segments of orbits provided that these segments are sparse enough in time. Systems with the specification property have very rich dynamics and the property has been used to investigate other properties like intrinsic ergodicity, topological mixing, positive topological entropy and others (see [5, 20]).
In order to study dynamical systems without Bowen’s specification, weaker notions were introduced, such as almost specification, weak specification and relative specification (we refer [16] for an overview).
The initial goal of this study was to answer a question from a recent work of Downarowicz and Weiss [7]. They asked about the relation between the weak specification property and a less explored variant of Bowen’s specification, called the vague specification property. The latter was introduced by Kamae [11]. We prove that the weak specification property implies the vague specification property and the converse does not hold. The conclusion follows from our main result, that establishes the equivalence of the asymptotic average shadowing and the vague specification properties, together with known relations between the weak specification and the asymptotic average shadowing properties.
More precisely, in [17], it is shown that both almost specification and weak specification properties imply the asymptotic average shadowing property. It is also known that there are systems with almost specification but without weak specification and vice-versa (see [14, 18]). As a consequence of the equivalence of the vague specification and the asymptotic average shadowing properties, one sees that the weak specification property implies the vague specification property and that the systems with the almost specification property but without the weak specification property serve as counterexamples for the converse. Therefore, the vague specification property can be considered one of the weakest forms of specification (see [16, 17]).
The equivalence of the vague specification and the asymptotic average shadowing properties also allows us to answer questions about the latter property. Indeed, we can use Kamae’s result (Proposition 5 in [11]), which shows that the vague specification property is inherited by factors, to address Question 10.6 in [14], confirming that the asymptotic average shadowing property is inherited by factors.
When studying the asymptotic average shadowing property, it is natural to ask its relation with the average shadowing property. It is already known that the asymptotic average shadowing property implies the average shadowing property. This result was first proved for surjective topological dynamical systems by Kulczycki, Kwietniak and Oprocha in [14]. Later, the result was improved by Wu, Oprocha and Chen in [22], who showed that the assumption of surjectivity could be dropped.
Roughly speaking, the dynamical Besicovitch pseudometric, defined on the phase space of a topological dynamical system, measures the asymptotic average distance between the iterates of two initial points. We show that for a surjective topological dynamical system $(X,T)$, the asymptotic average shadowing property is equivalent to $(X,T)$ being complete with respect to the dynamical Besicovitch pseudometric and having the average shadowing property. This provides a step toward answering Question 10.3 in [14].
Despite our efforts, we were unable to find any examples of systems that have the average shadowing property but are not Besicovitch complete. Thus, Question 10.3 in [14] remains unsolved and can be restated as: “Does the average shadowing property imply Besicovitch completeness?”
For surjective shift spaces, the average shadowing property is equivalent to the $\bar{d}$-shadowing property, introduced by Konieczny, Kupsa and Kwietniak in [13]. This equivalence was claimed without a proof in [13]. Since we rely on this equivalence in our work, we provide a proof of this statement for the sake of completeness.
We recall that the $\bar{d}$ and the Besicovitch pseudometrics are uniformly equivalent on shift spaces (see [17]). As a consequence, we obtain that, for surjective shift spaces, if we assume completeness with respect to the $\bar{d}$ pseudometric, then the $\bar{d}$-shadowing and the asymptotic average shadowing properties are equivalent.
In [6], the authors constructed a proximal shift space and a class of minimal shift spaces with the $\bar{d}$-shadowing property. The authors acknowledged Piotr Oprocha for the ideas that led to the construction of the minimal shift spaces. We prove that these examples are $\bar{d}$-complete, and therefore, as they are surjective, they have the asymptotic average shadowing property. This allows us to answer Question 10.2 in [14], by presenting minimal systems with the asymptotic average shadowing property.
Specification-like properties are the most common tools used to show that ergodic measures are entropy-dense in the space of invariant measures (see [16, 19, 20] for definitions and details). The proximal example and the class of minimal examples constructed in [6] are shown to have ergodic measures entropy-dense in the space of invariant measures. The surprise about such systems was that they do not seem to manifest any specification-like property, and therefore, different strategies are used to obtain the result. However, our results confirm that all these examples possess the vague specification property, which suggests that the vague specification property may imply entropy-density of ergodic measures.
2 Preliminaries
We write =$\mathbb{N}=\{1,2,\ldots\}$ and $\mathbb{N}_{0}=\mathbb{N}\cup\{0\}$. For every $A\subset\mathbb{N}_{0}$, we denote by $|A|\in\mathbb{N}_{0}\cup\{\infty\}$ the cardinality of $A$. For $n,m\in\mathbb{N}_{0}$ we write $[n,m)=\{i\in\mathbb{N}_{0}:n\leq i<m\}$, $(n,m]=\{i\in\mathbb{N}_{0}:n<i\leq m\}$ and $[n,m]=\{i\in\mathbb{N}_{0}:n\leq i\leq m\}$.
The upper asymptotic density of $A\subset\mathbb{N}_{0}$ is defined as
$\bar{d}(A)=\limsup_{n\to\infty}\frac{|A\cap[0,n)|}{n}.$
The lower asymptotic density of $A\subset\mathbb{N}_{0}$ is defined as
$\underline{d}(A)=\liminf_{n\to\infty}\frac{|A\cap[0,n)|}{n}.$
In case $\bar{d}(A)=\underline{d}(A)$, we define the asymptotic density of $A\subset\mathbb{N}_{0}$ by
$d(A)=\bar{d}(A)=\underline{d}(A).$
Throughout this work we assume that $X$ is a compact metrizable topological space and $\rho$ is a compatible metric on $X$. For simplicity, we assume that the diameter of $X$ with respect to $\rho$ is at most $1$. Our results do not depend on the choice of $\rho$.
We denote by $X^{\infty}$ the family of all $X$-valued sequences indexed by $\mathbb{N}_{0}$, that means,
$X^{\infty}=\left\{\left\{x_{n}\right\}_{n=0}^{\infty}:x_{n}\in X\text{ for all}n\in\mathbb{N}_{0}\right\}.$
Typically, we write $\underline{x}=\{x_{n}\}_{n=0}^{\infty}$ for elements of $X^{\infty}$. We denote by $S$ the shift operator acting on $X^{\infty}$, that is, $S(\underline{x})=\{x_{n+1}\}_{n=0}^{\infty}$. We always endow $X^{\infty}$ with the product topology, which is compatible with the metric
$\rho^{\infty}(\underline{x},\underline{z})=\sum_{n=0}^{\infty}\frac{\rho(x_{n},z_{n})}{2^{n+1}}.$
Note that the diameter of $X^{\infty}$ with respect to $\rho^{\infty}$ is again at most $1$. Immediately from the definition of $\rho^{\infty}$ we obtain the following result.
Proposition 2.1.
Let $\varepsilon>0$ and $k\in\mathbb{N}$. If $0<\delta<\frac{\varepsilon}{2^{k}}$ and $\rho^{\infty}(\underline{x},\underline{z})<\delta$ for $\underline{x},\underline{z}\in X$, then
$\max_{0\leq n<k}\rho(x_{n},z_{n})<\varepsilon.$
A pair $(X,T)$ is said to be a topological dynamical system if $T\colon X\to X$ is a continuous map. The orbit of a point $x\in X$ under $T$ is usually defined to be the set $\{T^{n}(x)\in X:n\in\mathbb{N}_{0}\}$ but, for our purposes, it is more convenient to treat orbits as sequences and we write $\underline{x}_{T}=\{T^{n}(x)\}_{n=0}^{\infty}\in X^{\infty}$. We write $\Gamma_{T}=\{\underline{x}_{T}\in X^{\infty}:x\in X\}$ for the space of all orbits under $T$. It is not hard to see that $\Gamma_{T}$ is a closed subspace of $X^{\infty}$, and therefore, $\Gamma_{T}$ is compact.
The Besicovitch pseudometric on $X^{\infty}$ is defined for $\underline{x},\underline{z}\in X^{\infty}$ as
$\rho_{B}(\underline{x},\underline{z})=\limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho(x_{i},z_{i}).$
Another pseudometric on $X^{\infty}$ is given for $\underline{x},\underline{z}\in X^{\infty}$ by
$\rho_{B}^{\prime}(\underline{x},\underline{z})$ $=\inf\left\{\varepsilon>0:\bar{d}(\{n\in\mathbb{N}_{0}:\rho(x_{n},z_{n})\geq\varepsilon\})<\varepsilon\right\}$
$=\inf\left\{\varepsilon>0:\underline{d}(\{n\in\mathbb{N}_{0}:\rho(x_{n},z_{n})<\varepsilon\})\geq 1-\varepsilon\right\}.$
Remark 2.2.
It is easy to see that the pseudometrics $\rho_{B}$ and $\rho_{B}^{\prime}$ on $X^{\infty}$ are uniformly equivalent (see Lemma 2 in [17] for a proof).
From the definition of $\rho_{B}^{\prime}$ we obtain the following equivalence.
Proposition 2.3.
Let $\underline{x},\underline{z}\in X^{\infty}$. Then $\rho_{B}^{\prime}(\underline{x},\underline{z})=0$ if and only if for every $\varepsilon>0$ we have
$d(\{n\in\mathbb{N}_{0}:\rho(x_{n},z_{n})<\varepsilon\})=1.$
Given a topological dynamical system $(X,T)$, the Besicovitch pseudometric on $X^{\infty}$ induces the dynamical Besicovitch pseudometric on $X$ which is defined for $x,z\in X$ as
$\rho^{T}_{B}(x,z)=\rho_{B}(\underline{x}_{T},\underline{z}_{T}).$
We define the notions of Cauchy sequence and completeness with respect to the dynamical Besicovitch pseudometric on $X$ in the usual way.
Definition 2.4.
Let $(X,T)$ be a topological dynamical system. We say that a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ is Besicovitch-Cauchy if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every $n,m\geq N$ it holds
$\rho^{T}_{B}(z^{(n)},z^{(m)})<\varepsilon.$
Definition 2.5.
We say that a topological dynamical system $(X,T)$ is Besicovitch complete if for every Besicovitch-Cauchy sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ there exists $z\in X$ (not necessarily unique) such that
$\lim_{n\to\infty}\rho^{T}_{B}(z^{(n)},z)=0.$
Remark 2.6.
The notion of Cauchy sequences and completeness for the Besicovitch pseudometric on $X^{\infty}$ could also be defined accordingly, and replicating the proof of Proposition 2 in [2] we see that $X^{\infty}$ is always Besicovitch complete.
3 Pseudo-orbits for $(X,T)$
Originally, pseudo-orbits were defined as sequences of points such that, starting from the second point in the sequence, every point is close to the image of the previous one. Roughly speaking, a pseudo-orbit is distinguished from an actual orbit only by small errors at each iteration.
Definition 3.1.
A sequence $\underline{x}\in X^{\infty}$ is called a $\delta$-pseudo-orbit for $(X,T)$ if $\rho(T(x_{n}),x_{n+1})<\delta$ for every $n\in\mathbb{N}_{0}.$
Allowing the errors to be small in different sense than the uniform one, like vanishing in the limit or on average, motivated definitions of different types of pseudo-orbits (we refer [9, 14, 19, 16] for more details). One of these generalized notions is the asymptotic pseudo-orbit.
Definition 3.2.
A sequence $\underline{x}\in X^{\infty}$ is called an asymptotic pseudo-orbit for $(X,T)$ if
$\lim\limits_{n\to\infty}\rho(T(x_{n}),x_{n+1})=0.$
Blank [3] analyzed pseudo-orbits, called average pseudo-orbits, in which the average of the errors becomes small in a long term.
Definition 3.3.
A sequence $\underline{x}\in X^{\infty}$ is said to be a $\delta$-average pseudo-orbit for $(X,T)$ if there exists $N\in\mathbb{N}$ such that for every $n\geq N$ and $k\in\mathbb{N}_{0}$ it holds
$\frac{1}{n}\sum_{i=0}^{n-1}\rho(T(x_{i+k}),x_{i+k+1})<\delta.$
From Definitions 3.2 and 3.3 it is not hard to see that asymptotic pseudo-orbits are $\delta$-average pseudo-orbits for any arbitrary $\delta>0$.
Proposition 3.4.
If $\underline{x}\in X^{\infty}$ is an asymptotic pseudo-orbit for $(X,T)$, then $\underline{x}$ is a $\delta$-average pseudo-orbit for $(X,T)$ for every $\delta>0$.
Later, Gu [9] studied the case when the average of the errors vanishes in the limit.
Definition 3.5.
A sequence $\underline{x}\in X^{\infty}$ is said to be an asymptotic average pseudo-orbit for $(X,T)$ if
$\limsup_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho(T(x_{i}),x_{i+1})=0.$
Combining Remark 2.2 and Proposition 2.3 we obtain the following result.
Proposition 3.6.
A sequence $\underline{x}\in X^{\infty}$ is an asymptotic average pseudo-orbit for $(X,T)$ if and only if for every $\varepsilon>0$ it holds
$d(\{n\in\mathbb{N}_{0}:\rho(T(x_{n}),x_{n+1})<\varepsilon\})=1.$
The combination of Lemmas 3.1 and 3.3 in [16] implies the following proposition.
Proposition 3.7.
Let $(X,T)$ be a surjective topological dynamical system. For every asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$ there exists an asymptotic pseudo-orbit $\underline{z}\in X^{\infty}$ such that
$d\left(\left\{n\in\mathbb{N}_{0}:x_{n}\neq z_{n}\right\}\right)=0.$
In particular, we have $\rho_{B}(\underline{x},\underline{z})=0$.
In order to define the vague specification property, Kamae [11] studied an unnamed family of sequences in $X^{\infty}$ that we will call vague pseudo-orbit.
Definition 3.8.
We call $\underline{x}\in X^{\infty}$ a vague pseudo-orbit for $(X,T)$ if for any open neighborhood ${\mathcal{U}}$ of $\Gamma_{T}$ in $X^{\infty}$ it holds
$d(\{n\in\mathbb{N}_{0}:S^{n}(\underline{x})\in{\mathcal{U}}\})=1.$
It turns out that asymptotic average pseudo-orbits and vague pseudo-orbits are the same sequences. In order to prove the equivalence between asymptotic average pseudo-orbits and vague pseudo-orbits we will need some lemmas.
The first lemma is a consequence of the thickness of subsets of $\mathbb{N}_{0}$ with asymptotic density $1$.
Lemma 3.9.
Let $G\subset\mathbb{N}_{0}$. If $d(G)=1$, then for every $k\in\mathbb{N}$ it holds
$d(\{n\in\mathbb{N}_{0}:[n,n+k)\subset G\})=1.$
Proof.
Let $G\subset\mathbb{N}_{0}$ be such that $d(G)=1$ and $k\in\mathbb{N}$. Set $\tilde{J}=\mathbb{N}_{0}\backslash G$. Note that $d(G)=1$ implies $d(\tilde{J})=0$, that is,
$\lim_{n\to\infty}\frac{|\tilde{J}\cap[0,n)|}{n}=0.$
Set
$J=\bigcup_{j\in\tilde{J}}[j-k,j].$
By construction, we have $|J\cap[0,n)|\leq k|\tilde{J}\cap[0,n)|+k$ for every $n\in\mathbb{N}$. Consequently,
$d(J)=\lim_{n\to\infty}\frac{|J\cap[0,n)|}{n}\leq\lim_{n\to\infty}\frac{k|\tilde{J}\cap[0,n)|+k}{n}=0.$
Lastly, observe that
$\{n\in\mathbb{N}_{0}:[n,n+k)\subset G\}=\mathbb{N}_{0}\backslash J.$
Therefore,
$d(\{n\in\mathbb{N}_{0}:[n,n+k)\subset G\})=1.\qed$
Straightforward from the definition of $\rho^{\infty}$ we have the following lemma.
Lemma 3.10.
Let $\varepsilon>0$. If $\frac{1}{2^{k}}<\varepsilon$ and $\underline{z}\in X^{\infty}$ satisfies $\max_{1\leq n\leq k}\rho(z_{n},T^{n}(z_{0}))<\varepsilon$, then $\rho^{\infty}(\underline{z},\underline{z_{0}}_{T})<\varepsilon$.
The last ingredient follows from the compactness of the space of orbits in the product topology. Informally, it says that for any open set ${\mathcal{U}}\subset X^{\infty}$ containing $\Gamma_{T}$, there exists a sufficiently large $k\in\mathbb{N}$ such that if the first $k$ terms of $\underline{z}\in X^{\infty}$ are sufficiently close to an orbit, then $\underline{z}$ belongs to ${\mathcal{U}}$.
Lemma 3.11.
Let ${\mathcal{U}}\subset X^{\infty}$ be an open set containing $\Gamma_{T}$. Then there exist $\varepsilon>0$ and $k\in\mathbb{N}$ such that
$\left\{\underline{z}\in X^{\infty}:\max_{1\leq n\leq k}\rho(z_{n},T^{n}(z_{0}))<\varepsilon\right\}\subset{\mathcal{U}}.$
Proof.
For $\varepsilon>0$ and $x\in X$, we denote
$B^{\infty}(x,\varepsilon)=\{\underline{y}\in X^{\infty}:\rho^{\infty}(\underline{y},\underline{x}_{T})<\varepsilon\}.$
By compactness of $\Gamma_{T}$, there exist $\varepsilon>0$ and $x^{(1)},\ldots,x^{(m)}\in X$ such that
$\Gamma_{T}\subset\bigcup_{i=1}^{m}B^{\infty}(x^{(i)},\varepsilon)\subset\bigcup_{i=1}^{m}B^{\infty}(x^{(i)},2\varepsilon)\subset{\mathcal{U}}.$
Let $k\in\mathbb{N}$ be such that $\frac{1}{2^{k}}<\varepsilon$. Take $\underline{z}\in X^{\infty}$ so that
$\max_{1\leq n\leq k}\rho(z_{n},T^{n}(z_{0}))<\varepsilon.$
By Lemma 3.10, we have $\rho^{\infty}(\underline{z},\underline{z_{0}}_{T})<\varepsilon.$ Note that $\underline{z_{0}}_{T}\in B^{\infty}(x^{(i)},\varepsilon)$ for some $i\in[1,m]$. Therefore, by the triangle inequality, $\underline{z}\in B^{\infty}(x^{(i)},2\varepsilon)\subset{\mathcal{U}}$. ∎
Theorem 3.12.
Let $(X,T)$ be a topological dynamical system. A sequence $\underline{x}\in X^{\infty}$ is an asymptotic average pseudo-orbit for $(X,T)$ if and only if $\underline{x}$ is a vague pseudo-orbit for $(X,T)$.
Proof.
Let $\varepsilon>0$ and $\underline{x}\in X^{\infty}$ be a vague pseudo-orbit. By Proposition 3.6, it suffices to show that
$d(\{n\in\mathbb{N}_{0}:\rho(T(x_{n}),x_{n+1})<\varepsilon\})=1.$ (1)
Using the uniform continuity of $T$, take $0<\delta<\frac{\varepsilon}{2}$ such that $\rho(T(w),T(v))<~{}\frac{\varepsilon}{2}$ for every $w,v\in X$ with $\rho(w,v)<\delta$.
Set ${\mathcal{U}}=\{\underline{y}\in X^{\infty}:\exists z\in X\text{ s.t. }\rho(y_{0},z)<\delta\text{ and }\rho(y_{1},T(z))<\delta\}$. Note that ${\mathcal{U}}$ is an open set containing $\Gamma_{T}$. Since $\underline{x}$ is a vague pseudo-orbit, by definition, we have
$d(\{n\in\mathbb{N}_{0}:S^{n}(\underline{x})\in{\mathcal{U}}\})=1.$
Observe that for each $n\in\mathbb{N}_{0}$ so that $S^{n}(\underline{x})\in{\mathcal{U}}$ there exists $z_{n}\in X$ such that $\rho(x_{n},z_{n})<\delta$ and $\rho(x_{n+1},T(z_{n}))<\delta$. Thus, by the triangle inequality, we obtain
$\rho(T(x_{n}),x_{n+1})\leq\rho(T(x_{n}),T(z_{n}))+\rho(x_{n+1},T(z_{n}))<\varepsilon.$
Therefore,
$\{n\in\mathbb{N}_{0}:S^{n}(\underline{x})\in{\mathcal{U}}\}\subset\{n\in\mathbb{N}_{0}:\rho(T(x_{n}),x_{n+1})<\varepsilon\},$
and consequently, equation (1) holds.
For the converse, let $\underline{x}\in X^{\infty}$ be an asymptotic average pseudo-orbit for $(X,T)$. Fix $\varepsilon>0$ and $k\in\mathbb{N}$. Set
${\mathcal{V}}=\left\{\underline{z}\in X^{\infty}:\max_{1\leq j\leq k}\rho(z_{j},T^{j}(z_{0}))<\varepsilon\right\}.$
By Lemma 3.11, to prove that $\underline{x}$ is a vague pseudo-orbit, it suffices to show that
$d(\{n\in\mathbb{N}_{0}:S^{n}(\underline{x})\in{\mathcal{V}}\})=1.$ (2)
Using uniform continuity of $T$, take $0<\delta<\frac{\varepsilon}{k}$ so that for any $w,v\in X$ with $\rho(w,v)<\delta$, we have $\rho(T^{j}(w),T^{j}(v))<\frac{\varepsilon}{k}$ for $j\in[1,k]$.
Set
$\tilde{G}=\{n\in\mathbb{N}_{0}:\rho(T(x_{n}),x_{n+1})<\delta\}.$
Since $\underline{x}$ is an asymptotic average pseudo-orbit, it follows from Proposition 3.6 that $d(\tilde{G})=1$. Note that, by the choice of $\delta$, for every $m\in\tilde{G}$ and $j\in[1,k]$ we have
$\rho\left(T^{j}(x_{m}),T^{j-1}(x_{m+1})\right)<\frac{\varepsilon}{k}.$ (3)
[…]
By the triangle inequality, for every $j\in[1,k]$ we obtain the following estimate (see Figure 1):
$\rho\left(T^{j}(x_{n}),x_{n+j}\right)\leq\sum_{i=0}^{j-1}\rho\left(T^{j-i}(x_{n+i}),T^{j-i-1}(x_{n+i+1})\right)<\varepsilon,$
Figure 1: illustration of the triangle inequality.
that is,
$\max_{1\leq j\leq k}\rho\left(T^{j}(x_{n}),x_{n+j}\right)<\varepsilon.$
Therefore, $G\subseteq\{n\in\mathbb{N}_{0}:S^{n}(\underline{x})\in{\mathcal{V}}\},$ showing that equation (2) holds. ∎
4 Tracing properties
In the same way that the original definition of pseudo-orbit motivated generalizations, the classical shadowing property inspired alternative types of tracing properties.
The average shadowing property introduced by Blank [3] is one of the alternative notions.
Definition 4.1.
We say that a topological dynamical system $(X,T)$ has the average shadowing property if for every $\varepsilon>0$ there exists $\delta>0$ such that for every $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ for $T$ there exists $z\in X$ such that
$\rho_{B}(\underline{z}_{T},\underline{x})<\varepsilon.$
We say that $\underline{x}$ is $\varepsilon$-traced on average by $z$.
In [14], Kulczycki, Kwietinak and Oprocha provided an equivalent formulation for the average shadowing property for the case where the topological dynamical system is chain mixing.
Definition 4.2.
We say that a topological dynamical system $(X,T)$ is chain mixing if for any $\delta>0$ and $x,y\in X$ there is $N\in\mathbb{N}$ such that for every $n\in\mathbb{N}$ satisfying $n\geq N$ there exists a sequence $\{x_{i}\}_{i=0}^{n}\subset X$ with $x_{0}=x$ and $x_{n}=y$ so that $\rho(T(x_{i}),x_{i+1})<\delta$ for all $i\in[0,n)$.
It is not hard to see that $\delta$-pseudo-orbits are also $\delta$-average pseudo-orbits. Roughly speaking, Kulczycki, Kwietniak and Oprocha showed that it is enough to assure that pseudo-orbits can be traced on average.
Theorem 4.3 (Theorem 3.6 in [14]).
Let $(X,T)$ be a topological dynamical system. If $(X,T)$ is chain mixing, then the following conditions are equivalent:
  1. 1.
    $(X,T)$ has the average shadowing property;
  2. 2.
    for every $\varepsilon>0$ there is $\delta>0$ such that for every $\delta$-pseudo-orbit $\underline{x}\in X^{\infty}$ there exists $z\in X$ so that $\rho_{B}(\underline{z}_{T},\underline{x})<\varepsilon.$
Inspired by the works of Blank [3], and Eirola, Nevanlinna and Pilyugin [8], Gu introduced the asymptotic average shadowing property in [9].
Definition 4.4.
We say that a topological dynamical system $(X,T)$ has the asymptotic average shadowing property if for every asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$ for $(X,T)$ there exists $z\in X$ such that
$\rho_{B}(\underline{z}_{T},\underline{x})=0.$
The average shadowing property is known to be a consequence of the asymptotic average shadowing property (see [14, 22]). One can see that the converse is true provided that $(X,T)$ is surjective and Besicovitch complete.
Theorem 4.5.
Let $(X,T)$ be a surjective topological dynamical system. Then $(X,T)$ has the asymptotic average shadowing property if and only if $(X,T)$ is Besicovitch complete and has the average shadowing property.
Proof.
By Corollary 4.2 in [22], the asymptotic average shadowing property implies the average shadowing property. Moreover, by Corollary 30 in [15], if $(X,T)$ has the asymptotic average shadowing property, then $(X,T)$ is Besicovitch complete.
Conversely, assume that $(X,T)$ is Besicovitch complete and has the average shadowing property. Let $\underline{x}\in X^{\infty}$ be an asymptotic average pseudo-orbit. By Proposition 3.7, there exists an asymptotic pseudo-orbit $\underline{y}\in X^{\infty}$ such that
$\rho_{B}(\underline{x},\underline{y})=0.$ (4)
It follows from Proposition 3.4 that $\underline{y}$ is a $\delta$-average pseudo-orbit for any $\delta>0$. Therefore, by the average shadowing property, for every $\varepsilon>0$ there exists a point $z^{(\varepsilon)}\in X$ such that $\rho_{B}(\underline{z}_{T}^{(\varepsilon)},\underline{y})<\varepsilon.$
Using the above observation we pick a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ such that $\rho_{B}(\underline{z}_{T}^{(n)},\underline{y})<2^{-n}$ for every $n\in\mathbb{N}$, and for $m>n$ we have $\rho^{T}_{B}(z^{(n)},z^{(m)})<2^{-n}.$
The Besicovitch completeness of $(X,T)$ implies that there exists $z\in X$ such that
$\lim_{n\to\infty}\rho^{T}_{B}(z^{(n)},z)=0.$
Moreover, by construction, we have
$\lim_{n\to\infty}\rho_{B}(\underline{z}_{T}^{(n)},\underline{y})=0.$
Therefore, $\rho_{B}(\underline{z}_{T},\underline{y})=0$, and by (4), we conclude that $\rho_{B}(\underline{z}_{T},\underline{x})=0$. ∎
We now focus on another tracing property, namely, the vague specification property, introduced by Kamae in [11].
Definition 4.6.
We say that a topological dynamical system $(X,T)$ has the vague specification property if for every vague pseudo-orbit $\underline{x}\in X^{\infty}$ for $(X,T)$ there is $z\in X$ such that for every $\varepsilon>0$ the following holds:
$d\left(\{n\in\mathbb{N}_{0}:\rho(T^{n}(z),x_{n})<\varepsilon\}\right)=1.$
If we combine Propositions 2.3 and 2.2, we see that the notions of tracing used in Definitions 4.4 and 4.6 are actually equivalent. As a consequence of Theorem 3.12, we obtain the equivalence between Definitions 4.4 and 4.6.
Theorem 4.7.
A topological dynamical system $(X,T)$ has the vague specification property if and only if $(X,T)$ has the asymptotic average shadowing property.
Proof.
We first assume that $(X,T)$ has the vague specification property. Fix an asymptotic average pseudo-orbit $\underline{x}\in X^{\infty}$. By Theorem 3.12, $\underline{x}$ is also a vague pseudo-orbit. Hence, by the vague specification property, there exists $z\in X$ such that for every $\varepsilon>0$ it holds
$d\left(\{n\in\mathbb{N}_{0}:\rho(T^{n}(z),x_{n})<\varepsilon\}\right)=1.$
Applying Lemma 2.3 followed by Lemma 2.2, we see that $\rho_{B}(\underline{z}_{T},\underline{x})=0$. Therefore, $(X,T)$ has the asymptotic average shadowing property.
Conversely, assume that $(X,T)$ has the asymptotic average shadowing property and take $\underline{x}\in X^{\infty}$ which is a vague pseudo-orbit. By Theorem 3.12, $\underline{x}\in X^{\infty}$ is also an asymptotic average pseudo-orbit for $(X,T)$. Since $(X,T)$ has the asymptotic average shadowing property, there exists $z\in X$ such that $\rho_{B}(\underline{z}_{T},\underline{x})=0$. Applying Lemmas 2.2 and 2.3, we obtain
$d\left(\{n\in\mathbb{N}_{0}:\rho(T^{n}(z),x_{n})<\varepsilon\}\right)=1.$
Hence, $(X,T)$ has the vague specification property. ∎
5 Tracing properties on shift spaces
Let $\mathscr{A}=\{0,1\}$ be endowed with the discrete topology. We call the set $\mathscr{A}$ an alphabet and each element of $\mathscr{A}$ we call a letter. By the full shift over $\mathscr{A}$ we mean the infinite Cartesian product $\mathscr{A}^{\infty}$ of $\mathscr{A}$ indexed by $\mathbb{N}_{0}$, that is,
$\mathscr{A}^{\infty}=\left\{\left\{x_{n}\right\}_{n=0}^{\infty}:x_{n}\in\mathscr{A}\text{ for all }n\in\mathbb{N}_{0}\right\}.$
We endow the full shift $\mathscr{A}^{\infty}$ with the product topology. By the Tychonoff Theorem, the space $\mathscr{A}^{\infty}$ is compact. A metric compatible with the product topology on $\mathscr{A}^{\infty}$ is given for $x=\{x_{n}\}_{n=0}^{\infty},y=\{y_{n}\}_{n=0}^{\infty}\in\mathscr{A}^{\infty}$ by
$\rho(x,y)=\begin{cases}0,&\mbox{if }x=y,\\ 2^{-\min\{n\in\mathbb{N}_{0}:x_{n}\neq y_{n}\}},&\mbox{otherwise}.\end{cases}$
The shift map $\sigma\colon\mathscr{A}^{\infty}\to\mathscr{A}^{\infty}$ is defined by $\sigma(x)=\{x_{n+1}\}_{n=0}^{\infty}$ where $x=\{x_{n}\}_{n=0}^{\infty}\in\mathscr{A}^{\infty}$. A nonempty, closed, and $\sigma$-invariant set $X\subseteq\mathscr{A}^{\infty}$ is called a shift space over $\mathscr{A}$. When referring to a shift space and its properties, we will be implicitly assuming that we are dealing with the topological dynamical system $(X,\sigma)$. For aesthetic reasons, along this section we will adopt an upper index for sequences of points in $X$, that is, we write
$X^{\infty}=\{\{x^{(n)}\}_{n=0}^{\infty}:x^{(n)}\in X\text{ for all }n\in\mathbb{N}_{0}\}.$
We call a finite sequence of letters from $\mathscr{A}$ a word over $\mathscr{A}$. The number of letters in a word $w$ is called the length of $w$. We write $|w|$ for the length of a word $w$. The only sequence with length 00 is called the empty word and is denoted by $\lambda$.
Let $x=\{x_{n}\}_{n=0}^{\infty}\in\mathscr{A}^{\infty}$ and $i,j\in\mathbb{N}_{0}$. For $i\leq j$, we write $x_{[i,j)}$ for the word $x_{i}x_{i+1}\ldots x_{j-1}$ over $\mathscr{A}$ of length $j-i$ and we agree that $x_{[i,i)}=\lambda$. We will also write $x_{[k,\infty)}$ to denote $\sigma^{k}(x)$ for $k\in\mathbb{N}_{0}$.
We write $\mathscr{A}^{*}$ for the set of all words over $\mathscr{A}$. Each word $w\in\mathscr{A}^{*}$ determines the cylinder set $[w]=\{x\in\mathscr{A}^{\infty}:x_{[0,|w|)}=w\}$. The language of a shift space $X$ is $\mathcal{L}(X)=\left\{w\in\mathscr{A}^{*}:[w]\cap X\neq\emptyset\right\}$.
For every $n\in\mathbb{N}_{0}$, we denote by $\mathcal{L}_{n}(X)\subset\mathscr{A}^{*}$ the set of all words $w\in\mathcal{L}(X)$ such that $|w|=n$. Hence, it holds
$\mathcal{L}(X)=\bigcup_{n\in\mathbb{N}_{0}}\mathcal{L}_{n}(X).$
By the concatenation of the words $u=u_{1}\ldots u_{k}\in\mathscr{A}^{*}$ and $v=v_{1}\ldots v_{m}\in\mathscr{A}^{*}$ we mean the word $u_{1}\ldots u_{k}v_{1}\ldots v_{m}$, and we denote it by $uv$. We say that a word $u\in\mathscr{A}^{*}$ is a prefix (resp. suffix) of a word $w\in\mathscr{A}^{*}$ if there exists $v\in\mathscr{A}^{*}$ such that $uv=w$ (resp. $vu=w$). For $n\in\mathbb{N}$, the word $0^{n}=0\ldots 0\in\mathscr{A}^{*}$ (resp. $1^{n}=1\ldots 1$) stands for the word of length $n$ composed exclusively of 00’s (resp. $1$’s). If $w\in\mathscr{A}^{*}$, then $w^{\infty}=www\ldots\in\mathscr{A}^{\infty}$ represents the infinite self-concatenation of $w$.
We say that a shift space $X$ is mixing if for every pair of words $u,v\in\mathcal{L}(X)$ there exists $N\in\mathbb{N}$ such that for every $n\geq N$ there is a word $w\in\mathcal{L}_{n}(X)$ such that $uwv\in\mathcal{L}(X)$.
Let $\mathcal{B}\subset\mathscr{A}^{*}$ be a collection of words. We say that $\mathcal{B}$ is a code for a shift space $X$, if $X$ is the collection of all possible infinite concatenations of words from ${\mathcal{B}}$ and their shifts. Moreover, $X$ is called coded shift in this case.
We denote an oriented $\mathscr{A}$-labelled (multi) graph by a triple $G=\left(V,E,\tau\right)$, where $V$ is a set of vertices, $E$ is a set of edges, and a label map $\tau\colon E\to\mathscr{A}$. Every $e\in E$ has two end points, an initial vertex $i(e)\in V$ and terminal vertex $t(e)\in V$. Let $\ell\in\mathbb{N}$. A sequence of edges $\{e_{1},\ldots,e_{\ell}\}\subset E$ is called a path of length $\ell$ if there exists a sequence of vertices $\{v_{1},\ldots,v_{\ell}\}\subset V$ with $i(e_{i})=v_{i}$, and $t(e_{i})=v_{i+1}$ for every $1\leq i<\ell$.
We define the shift $X_{G}$ to be the set of all sequences $x\in\mathscr{A}^{\infty}$ such that $x$ is the sequence of labels of an infinite path on a graph $G$ and their shifts. A shift space $X$ is called sofic if there exists a labelled graph $G=\left(V,E,\tau\right)$ such that the set of vertices $V$ is finite and $X=X_{G}$.
The hereditary closure of a shift space $X$ is
[…]
A shift space $X$ is called hereditary if $\tilde{X}=X$.
We say that $w\in\mathcal{L}(X)$ is a synchronizing word for a shift space $X$ if $uw,wv\in\mathcal{L}(X)$ implies $uwv\in\mathcal{L}(X)$. A shift space $X$ is called synchronized if it is transitive and has a synchronizing word. In the particular case of hereditary shifts, it is not hard to see that we can assume the synchronizing word to be composed only by 00’s. More precisely, we have the following lemma whose proof is left to the reader.
Lemma 5.1.
Let $X$ be a hereditary shift space. If $w\in\mathcal{L}_{n}(X)$ is a synchronizing word for $X$, then $0^{n}\in\mathcal{L}(X)$ is also a synchronizing word for $X$.
As a consequence of Lemma 5.1, we see that if a surjective shift space $X$ is hereditary and synchronized, then there exists a fixed integer $n\in\mathbb{N}$ such that we can connect any word $u\in\mathcal{L}(X)$ with any point $x\in X$ using a word $0^{n}$.
Lemma 5.2.
Let $X$ be a hereditary surjective shift space and $w\in~{}\mathcal{L}_{n}(X)$ be a synchronizing word for $X$. If $u\in\mathcal{L}(X)$ and $x\in X$, then $u0^{n}x\in X$.
Proof.
Note that, by hereditarity, $u0^{m}\in\mathcal{L}(X)$ for every $u\in\mathcal{L}(X)$ and $m\in\mathbb{N}$.
Since $X$ is surjective, for every $x\in X$ and $n\in\mathbb{N}$ there exists a word $v\in\mathcal{L}_{n}(X)$ such that $vx\in X$. By hereditarity, we can assume that $v=0^{n}$, that is, $0^{n}x\in X$.
It follows from Lemma 5.1 that $0^{n}$ is a synchronizing word for $X$. Hence, for every $v\in\mathcal{L}(X)$ that is prefix of $x$ we have $u0^{n}v\in\mathcal{L}(X)$, and consequently, $u0^{n}x\in X$. ∎
The original specification property introduced by Bowen [4] for general topological dynamical systems has an equivalent formulation for shift spaces.
Definition 5.3.
Let $X$ be a shift space. We say that $X$ has the specification property if there exists $n\in\mathbb{N}$ such that for every $u,v\in\mathcal{L}(X)$ there is $w\in\mathcal{L}_{n}(X)$ such that $uwv\in\mathcal{L}(X)$.
The class of shift spaces with the specification property is quite broad and contains, for example, the class of mixing sofic shifts.
Proposition 5.4 (Section 6 of [21]).
Every mixing sofic shift space has the specification property.
Remark 5.5.
We remark that there are mixing shift spaces without the specification property (see [10, 16]).
In 1988, Bertrand [1] proved that shift spaces with specification always have synchronizing words.
Theorem 5.6 (Theorem 1 in [1]).
Let $X$ be a shift space. If $X$ has the specification property, then $X$ is synchronized.
We will abuse the notation and denote by $\bar{d}$ also the pseudometric on $\mathscr{A}^{\infty}$ induced by the upper asymptotic density given for $x,y\in\mathscr{A}^{\infty}$ by
$\bar{d}(x,y)=\bar{d}(\left\{n\in\mathbb{N}_{0}:x_{n}\neq y_{n}\right\})=\limsup_{n\to\infty}\frac{1}{n}\left|\left\{i\in[0,n):x_{i}\neq y_{i}\right\}\right|.$
Proposition 5.7 (Corollary 5 in [17]).
The dynamical Besicovitch and the $\bar{d}$ pseudometrics on $\mathscr{A}^{\infty}$ are uniformly equivalent.
We will discuss convergence of sequences of points in $\mathscr{A}^{\infty}$ with respect to the $\bar{d}$ pseudometric. We remark that since $\bar{d}$ is a pseudometric, a limit point, if exists, might not be unique. We define the notion of Cauchy sequence and completeness with respect to the $\bar{d}$ pseudometric in the natural way.
Definition 5.8.
We say that a sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ is $\bar{d}$-Cauchy if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every $n,m\geq N$ it holds
$\bar{d}(z^{(n)},z^{(m)})<\varepsilon.$
Definition 5.9.
We say that a shift space $X$ is $\bar{d}$-complete if for every $\bar{d}$-Cauchy sequence $\{z^{(n)}\}_{n=1}^{\infty}\subset X$ there exists $z\in X$ such that
$\lim_{n\to\infty}\bar{d}(z^{(n)},z)=0.$
By the uniform equivalence between the Besicovitch and $\bar{d}$ pseudometrics (Proposition 5.7), it follows from Corollary 30 in [15] that the asymptotic average shadowing property is a sufficient condition for $\bar{d}$-completeness.
Proposition 5.10.
Let $X$ be a shift space. If $X$ has the asymptotic average shadowing property, then $X$ is $\bar{d}$-complete.
In the particular case of shift spaces, a notion of tracing called $\bar{d}$-shadowing property, which is closely related to the average shadowing property, was introduced by Konieczny, Kupsa and Kwietniak in [13].
Definition 5.11.
We say that a shift space $X$ has the $\bar{d}$-shadowing property if for every $\varepsilon>0$ there exists $N\in\mathbb{N}$ such that for every sequence $\{w^{(i)}\}_{i=1}^{\infty}\subset{\mathcal{L}}(X)$ with $|w^{(i)}|\geq N$ for every $i\in\mathbb{N}$ there exists $x\in X$ such that
$\bar{d}(x,w)<\varepsilon,$
where $w=w^{(1)}w^{(2)}\ldots\in\mathscr{A}^{\infty}$.
Remark 5.12.
We remark that given a sequence of words $\{w^{(i)}\}_{i=1}^{\infty}\subset{\mathcal{L}}(X)$ their concatenation $w=w^{(1)}w^{(2)}\ldots$ might not belong to $X$.
As mentioned in [13], a shift space with the average shadowing property also possesses the $\bar{d}$-shadowing property. Moreover, if we assume surjectivity of the shift space, then the $\bar{d}$-shadowing property implies the average shadowing property. Due to the lack of proofs for these statements, for the sake of completeness, we will prove them.
Lemma 5.13.
For every $\delta>0$ there exists $N\in\mathbb{N}$ such that if a sequence of words $\{w^{(i)}\}_{i=1}^{\infty}\subset\mathcal{L}(X)$ satisfies $|w^{(i)}|\geq N$ for every $i\in\mathbb{N}$, then there exists a $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ such that
$\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta,$
where $w=w^{(1)}w^{(2)}\ldots\in\mathscr{A}^{\infty}$.
Proof.
Let $\delta>0$. Take $m\in\mathbb{N}$ so that $2^{-m}<\frac{\delta}{2}$ and $N\in\mathbb{N}$ satisfying $\frac{m}{N}<\frac{\delta}{2}$. Fix a sequence of words $\{w^{(i)}\}_{i=1}^{\infty}\subset\mathcal{L}(X)$ with $|w^{(i)}|\geq N$ for every $i\in\mathbb{N}$. Let $w=w^{(1)}w^{(2)}\ldots\in\mathscr{A}^{\infty}$ denote the concatenation of these words.
Set $n_{1}=0$ and $n_{i}=|w^{(1)}\ldots w^{(i-1)}|$ for $i\geq 2$. We now construct the sequence $\underline{x}=\{x^{(n)}\}_{n=0}^{\infty}\in X^{\infty}$ as follows. For each $n\in\mathbb{N}$ there exists $i=i(n)\in\mathbb{N}$ such that $n\in[n_{i},n_{i+1})$. Hence, we pick
$x^{(n)}\in\left[w^{(i)}_{[n-n_{i},|w^{(i)}|)}\right]\cap X.$
Set $E=\bigcup_{i\geq 2}(n_{i}-m,n_{i}]$. Note that if $n\in\mathbb{N}_{0}\backslash E$, then for its correspondent $i$ as above, it holds
$\left|w^{(i)}_{[n-n_{i},|w^{(i)}|)}\right|>m.$
Thus, by construction, for every $n\in\mathbb{N}_{0}\backslash E$ we have $\rho(\sigma^{n}(w),x^{(n)})\leq 2^{-m}$. We observe that $\bar{d}(E)\leq\frac{m}{N}$, which implies $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\frac{\delta}{2}$.
It remains to prove that $\underline{x}\in X^{\infty}$ is a $\delta$-average pseudo-orbit. Note that, also by construction, for every $n\in\mathbb{N}_{0}\backslash E$ it holds $x^{(n)}_{[1,m]}=x^{(n+1)}_{[0,m)}$, that is,
$\rho(\sigma(x^{(n)}),x^{(n+1)})\leq 2^{-m}.$ (5)
Fix $n\geq N$ and $k\in\mathbb{N}_{0}$. We first find an upper bound for $|[k,k+n)\cap E|$. Observe that in the worst case scenario we have $n_{i+1}-n_{i}=N$ for all $i\in\mathbb{N}$. Thus,
$|[k,k+n)\cap E|\leq\frac{nm}{N}<\frac{n\delta}{2}.$ (6)
We can split the sum below into terms in $E$ and outside $E$. Since the diameter of $X$ is assumed to be at most $1$, using (5) and (6), we obtain
$\frac{1}{n}\sum_{i=0}^{n-1}\rho(\sigma(x^{(k+i)}),x^{(k+i+1)})$ $\leq\frac{\left(n-|[k,k+n\right)\cap E|)\cdot 2^{-m}}{n}+\frac{|[k,k+n)\cap E|}{n}$
$<2^{-m}+\frac{\delta}{2}<\delta.\qed$
Theorem 5.14.
Let $X$ be a shift space. If $X$ has the average shadowing property, then $X$ has the $\bar{d}$-shadowing property.
Proof.
Let $\varepsilon>0$. By the uniform equivalence between the Besicovitch and the $\bar{d}$ pseudometrics, there exists $0<\varepsilon^{\prime}<\varepsilon$ such that for every $w,z\in\mathscr{A}^{\infty}$ satisfying $\rho_{B}(\underline{w}_{\sigma},\underline{z}_{\sigma})<\varepsilon^{\prime}$, we have $\bar{d}(w,z)<\varepsilon$.
Let $0<\delta\leq\frac{\varepsilon^{\prime}}{2}$ be such that every $\delta$-average pseudo-orbit is $\frac{\varepsilon^{\prime}}{2}$-traced on average by a point in $X$ and $N\in\mathbb{N}$ be provided by Lemma 5.13. Fix a sequence of words $\{w^{(i)}\}_{i=1}^{\infty}\subset\mathcal{L}(X)$ satisfying $|w^{(i)}|\geq N$ for every $i\in\mathbb{N}$.
By Lemma 5.13, there exists a $\delta$-average pseudo-orbit $\underline{x}\in X^{\infty}$ so that $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta$. The average shadowing property yields the existence $z\in X$ such that $\rho_{B}(\underline{x},\underline{z}_{\sigma})<\frac{\varepsilon^{\prime}}{2}$. Therefore, by the triangle inequality, we obtain $\rho_{B}(\underline{w}_{\sigma},\underline{z}_{\sigma})<\varepsilon^{\prime}$. Finally, by the choice of $\varepsilon^{\prime}$, we conclude that $\bar{d}(w,z)<\varepsilon$. ∎
Our proof for a converse of Theorem 5.14 requires the shift space to be chain mixing. Lemma 9 in [13] guarantees that, in our context, surjectivity is sufficient.
Lemma 5.15 (Lemma 9 in [13]).
Let $X$ be a surjective shift space. If $X$ has the $\bar{d}$-shadowing property, then $X$ is chain mixing.
Theorem 5.16.
Let $X$ be a surjective shift space. If $X$ has the $\bar{d}$-shadowing property, then $X$ has the average shadowing property.
Proof.
Let $\varepsilon>0$. By the uniform equivalence between the Besicovitch and the $\bar{d}$ pseudometrics, there exists $\varepsilon^{\prime}>0$ such that $\rho_{B}(\underline{w}_{\sigma},\underline{z}_{\sigma})<\frac{\varepsilon}{2}$ for any $w,z\in\mathscr{A}^{\infty}$ so that $\bar{d}(w,z)<\varepsilon^{\prime}$.
Let $N\in\mathbb{N}$ be provided by the $\bar{d}$-shadowing property for $\varepsilon^{\prime}$. Without loss of generality, we may assume that $2^{-N}<\frac{\varepsilon}{2}$. By Lemma 5.15, $X$ is chain-mixing, so by Theorem 4.3, it remains to prove that there exists $\delta>0$ such that every $\delta$-pseudo-orbit is $\varepsilon$-traced on average by some point $z\in X$.
Let $m\in\mathbb{N}$ be such that $\frac{N}{m}<\frac{\varepsilon}{2}$ and $0<\delta<\frac{\varepsilon}{2}$ be such that $\delta\leq 2^{-m}$. Fix a $\delta$-pseudo-orbit $\underline{x}\in X^{\infty}$. Note that $\delta\leq 2^{-m}$ implies $x^{(n)}_{[1,m]}=x^{(n+1)}_{[0,m)}$ for every $n\in\mathbb{N}_{0}$. As a consequence, for every $k\in[0,m)$ it holds
$x^{(n)}_{[k,m)}=x^{(n+k)}_{[0,m-k)}.$ (7)
For each $i\in\mathbb{N}$ we set $w^{(i)}=x^{(i-1)m}_{[0,m)}\in\mathcal{L}_{m}(X)$. Let $w=w^{(1)}w^{(2)}\ldots\in\mathscr{A}^{\infty}$ denote the concatenation of $w^{(i)}$’s.
For every $n\in\mathbb{N}_{0}$ there exists $i=i(n)\in\mathbb{N}$ such that $n\in[(i-1)m,im)$. It follows from (7) that
$w_{[n,im)}=x^{(i-1)m}_{[n-(i-1)m,m)}=x^{(n)}_{[0,im-n)}.$ (8)
We will use an argument similar to the one in Lemma 5.13 to prove that $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\delta$. Set $E=\cup_{i\in\mathbb{N}}(im-N,im]$. It follows from (8) that for every $n\in\mathbb{N}_{0}\backslash E$ it holds $\rho(\sigma^{n}(w),x^{(n)})\leq 2^{-N}$. Note that $\bar{d}(E)\leq\frac{N}{m}$, thus $\rho_{B}(\underline{w}_{\sigma},\underline{x})<\frac{\varepsilon}{2}.$
Now we use the $\bar{d}$-shadowing property to find $z\in X$ such that $\bar{d}(w,z)<\varepsilon^{\prime}$. Lastly, by the triangle inequality, the choice of $\varepsilon^{\prime}$ implies $\rho_{B}(\underline{x},\underline{z}_{\sigma})<\varepsilon.$
Since the $\bar{d}$-shadowing and the average shadowing properties are equivalent for surjective shift spaces, the equivalence of $\bar{d}$ and Besicovitch pseudometrics, together with Theorem 4.5, yields through the following corollary.
Corollary 5.17.
Let $X$ be a surjective shift space. Then $X$ has the asymptotic average shadowing property if and only if $X$ is $\bar{d}$-complete and has the $\bar{d}$-shadowing property.
6 Examples of shift spaces with the vague specification property
6.1 An example of proximal shift space
We will show that the shift space $Z$, defined in Example 21 in [6], has the vague specification property. Since it is known that $Z$ has the $\bar{d}$-shadowing property, it is enough to show that $Z$ is $\bar{d}$-complete and use Corollary 5.17. Before showing that $Z$ is $\bar{d}$-complete we recall its construction following [6].
Example 6.1.
Let $n\in\mathbb{N}$. We construct a sofic shift $Z_{n}$ represented by an oriented labeled graph $G_{n}=(V_{n},E_{n},\tau_{n})$ with vertex set $V_{n}=\{v_{0},v_{1},\ldots,v_{10^{n}-1}\}$ whose edge set $E_{n}$ and labels $\tau_{n}:E_{n}\to\{0,1\}$ are given by
  • For every $0\leq k<10^{n}$ there is an edge from $v_{k}$ to $v_{k+1\mod 10^{n}}$ with label 00.
  • For every $1\leq k\leq 10^{n}-2^{n}$ there is an edge from $v_{k}$ to $v_{k+1}$ with label $1$.
  • There is an edge from $v_{10^{n}-2^{n}}$ to $v_{10^{n}-2^{n}+2}$ with label 00.
Let $Z=\bigcap_{n=1}^{\infty}Z_{n}$. We remark that for every $n\in\mathbb{N}$ we have $Z_{n+1}\not\subset Z_{n}$ and $Z_{n}\not\subset Z_{n+1}$. For each $n\in\mathbb{N}$ we set $Y_{n}=\bigcap_{j=1}^{n}Z_{n}$. This defines a decreasing sequence of mixing sofic shifts $Y_{1}\supseteq Y_{2}\supseteq\dots$ with $Z=\bigcap_{n=1}^{\infty}Y_{n}$.
We denote
$\tilde{E}_{j}=\{n\in\mathbb{N}:n\geq 10^{j}-2^{j}\mod 10^{j}\}\quad\text{and}\quad E_{i}=\bigcup_{j\geq i}\tilde{E}_{j}.$ (9)
Note that
$\bar{d}(E_{i})\leq\frac{1}{4\cdot 5^{i-1}}<5^{-i+1}.$ (10)
The crucial observation about $Z$, that we will repeatedly use below, is that if $n\in\mathbb{N}$ and $y\in Y_{n}$, then it is enough to set $y_{j}$ to 00 for every $j\in E_{n}$ to obtain a point in $Z$.
Theorem 6.2.
The shift space $Z$ is $\bar{d}$-complete.
Proof.
We let $\tilde{E}_{j}$ and $E_{i}$ to be defined as in (9). Let $\{x^{(n)}\}_{n=1}^{\infty}$ be a $\bar{d}$-Cauchy sequence in $Z$. Passing to a subsequence, we can assume that for every $n\in\mathbb{N}$ it holds
$\bar{d}(x^{(n)},x^{(n+1)})<5^{-n-1}.$ (11)
By the triangle inequality and (11), for every $m\in\mathbb{N}$ with $m\geq n$, we have
$\bar{d}(x^{(n)},x^{(m)})<5^{-n}.$ (12)
We will construct a strictly increasing sequence of integers $\{\ell_{n}\}_{n=2}^{\infty}\subset\mathbb{N}$ and a sequence of points $\{z^{(n)}\}_{n=1}^{\infty}\subset Z$ so that for every $n\in\mathbb{N}$ the word $z^{(n)}_{[0,\ell_{n+1})}$ is a prefix of $z^{(n+1)}$, that is,
$z^{(n)}_{[0,\ell_{n+1})}=z^{(n+1)}_{[0,\ell_{n+1})}.$ (13)
Using (13) we define $x\in Z$ to be the unique point satisfying $x_{[0,\ell_{n})}=z^{(n)}_{[0,\ell_{n})}$ for every $n\geq 1$. We are going to show that $x$ is a $\bar{d}$-limit of $\{x^{(n)}\}_{n=1}^{\infty}$. Equivalently, $x$ is given by
$x=z^{(1)}_{[0,\ell_{2})}z^{(2)}_{[\ell_{2},\ell_{3})}\cdots.$
Note that $Y_{n}$ is a hereditary shift with specification for every $n\in\mathbb{N}$. Therefore, by Lemmas 5.1 and 5.2, for every $n\in\mathbb{N}$ there exists $m_{n}\in\mathbb{N}$ such that $u_{n}=0^{m_{n}}$ is a synchronizing word for $Y_{n}$.
We construct an auxiliary sequence of points $\{y^{(n)}\}_{n=1}^{\infty}$ as
$y^{(n)}_{i}=\begin{cases}0,\quad&\text{ if }i\in E_{n},\\ x^{(n)}_{i},&\text{ otherwise.}\end{cases}$
Note that the definition of $Z$ implies that $y^{(n)}\in Z$. Observe that, using (10), for every $n\in\mathbb{N}$ we have
$\bar{d}(y^{(n)},x^{(n)})\leq\bar{d}(E_{n})<5^{-n+1}.$ (14)
Let $\ell_{2}\in\mathbb{N}$ be such that
  1. 1.
    $\frac{|u_{2}|}{\ell_{2}}<5^{-2}$;
  2. 2.
    |$\left|\left\{i\in[0,\ell):y^{(2)}_{i}\neq x^{(2)}_{i}\right\}\right|<\ell\cdot 5^{-1}$ for every $\ell\geq\ell_{2}$ (possible by (14)).
Set $z^{(1)}=y^{(1)}$ and
$z^{(2)}=y^{(1)}_{[0,\ell_{2})}u_{2}y^{(2)}_{[\ell_{2}+|u_{2}|,\infty)}\in Y_{2}=Z_{1}\cap Z_{2}.$
We have $\bar{d}(z^{(2)},y^{(2)})=0$. To see that $z^{(2)}\in Z$, it is enough to note that $z^{(2)}_{i}=0$ for every $i\in E_{2}$ and $z^{(2)}\in Y_{2}$.
We proceed with the construction of $z^{(n)}\in Z$ inductively as follows. For $n>2$ we will take $\ell_{n}\in\mathbb{N}$ satisfying the following conditions:
  1. (1)
    $\ell_{n}\geq 2\ell_{n-1}$;
  2. (2)
    $\frac{|u_{n}|}{\ell_{n}}<5^{-n}$;
  3. (3)
    |$\left|\left\{i\in[0,\ell):y^{(n)}_{i}\neq x^{(n)}_{i}\right\}\right|<\ell\cdot 5^{-n+1}$ for every $\ell\geq\ell_{n}$ (possible by (14));
  4. (4)
    |$\left|\left\{i\in[0,\ell):x^{(n)}_{i}\neq x^{(m)}_{i}\right\}\right|<\ell\cdot 5^{-n}$ for every $\ell\geq\ell_{n}$ (possible by (12));
  5. (5)
    $|E_{n}\cap[0,\ell)|<\ell\cdot 5^{-n+1}$ for every $\ell\geq\ell_{n}$ (possible by (10)).
We define $z^{(n)}$ by
$z^{(n)}=z^{(n-1)}_{[0,\ell_{n})}u_{n}y^{(n)}_{[\ell_{n}+|u_{n}|,\infty)}\in Y_{n}=\bigcap_{j=1}^{n}Z_{j}.$
For every $n\in\mathbb{N}$ we have $\bar{d}(z^{(n)},y^{(n)})=0$. To see that $z^{(n)}\in Z$, it is enough to note that $z^{(n)}\in Y_{n}$ and $z^{(n)}_{i}=0$ for every $i\in E_{n}$. Also, for every $n\in\mathbb{N}$ we have
|$\left|\left\{i\in[\ell_{n},\ell_{n+1}):z^{(n)}_{i}\neq y^{(n)}_{i}\right\}\right|\leq|u_{n}|.$ (15)
By the triangle inequality, it follows from condition (3) that our choice of $\ell_{n}$, together with (14) and (15), guarantees
|$\begin{split}\left|\left\{i\in[\ell_{n},\ell_{n+1}):x_{i}\neq x^{(n)}_{i}\right\}\right|&=\left|\left\{i\in[\ell_{n},\ell_{n+1}):z^{(n)}_{i}\neq x^{(n)}_{i}\right\}\right|\\ &\leq\left|\left\{i\in[\ell_{n},\ell_{n+1}):y^{(n)}_{i}\neq x^{(n)}_{i}\right\}\right|+|u_{n}|\\ &\leq\ell_{n+1}\cdot 5^{-n+1}+|u_{n}|.\end{split}$ (16)
Conditions (1) and (2) mean that our choice of $\ell_{n}$ guarantees that for every $n,k\in\mathbb{N}$ such that $k>n$ we have
$\sum_{j=n}^{k-1}|u_{j}|\leq\sum_{j=n}^{k-1}\ell_{j}\cdot 5^{-j}<2\ell_{k-1}\cdot 5^{-n}\leq\ell_{k}\cdot 5^{-n}.$ (17)
It remains to show that $\lim_{n\to\infty}\bar{d}(x^{(n)},x)=0$. Fix $n\in\mathbb{N}$. For every $\ell\geq\ell_{n+1}$ there exists $k\geq n+1$ such that $\ell\in[\ell_{k},\ell_{k+1})$. Note that
|$\left|\left\{i\in[0,\ell):x_{i}\neq x^{(n)}_{i}\right\}\right|\leq\left|\left\{i\in[0,\ell_{n+1}):x_{i}\neq x^{(n)}_{i}\right\}\right|+\left|\left\{i\in[\ell_{n+1},\ell):x_{i}\neq x^{(n)}_{i}\right\}\right|.$
By the triangle inequality, we have
|$\left|\left\{i\in[\ell_{n+1},\ell):x_{i}\neq x^{(n)}_{i}\right\}\right|\leq$ $\sum_{j=n+1}^{k-1}\left|\left\{i\in[\ell_{j},\ell_{j+1}):x^{(n)}_{i}\neq x^{(j)}_{i}\right\}\right|$
$+\sum_{j=n+1}^{k-1}\left|\left\{i\in[\ell_{j},\ell_{j+1}):x_{i}\neq x^{(j)}_{i}\right\}\right|$
+|$+\left|\left\{i\in[\ell_{k},\ell):x^{(n)}_{i}\neq x^{(k)}_{i}\right\}\right|+\left|\left\{i\in[\ell_{k},\ell):x_{i}\neq x^{(k)}_{i}\right\}\right|.$
By condition (4) for the choice of $\ell_{n}$’s and inequality (16), we obtain
|$\begin{split}\left|\left\{i\in[0,\ell):x_{i}\neq x^{(n)}_{i}\right\}\right|\leq&\ell_{n+1}+\sum_{j=n+1}^{k-1}\ell_{j+1}\cdot 5^{-n}+\sum_{j=n+1}^{k-1}\left(\ell_{j+1}\cdot 5^{-j+1}+|u_{j}|\right)\\ &+\ell\cdot 5^{-n}+|u_{k}|+|E_{k}\cap[\ell_{k},\ell)|.\end{split}$ (18)
Note that condition (1) implies
$\sum_{j=n+1}^{k-1}\ell_{j+1}\cdot 5^{-j+1}\leq\sum_{j=n+1}^{k-1}\ell_{j+1}\cdot 5^{-n}\leq 2\ell_{k}\cdot 5^{-n}.$
As a consequence, from (17) we obtain
$\begin{split}\sum_{j=n+1}^{k-1}\left(\ell_{j+1}\cdot 5^{-j+1}+|u_{j}|\right)&\leq\sum_{j=n+1}^{k-1}|u_{j}|+\sum_{j=n+1}^{k-1}\ell_{j+1}\cdot 5^{-j+1}\\ &\leq\ell_{k}\cdot 5^{-n-1}+2\ell_{k}\cdot 5^{-n}\\ &\leq 3\ell_{k}\cdot 5^{-n}.\end{split}$ (19)
Therefore, using conditions (1) and (4) together with (19), it follows from (18) that
|$\begin{split}\left|\left\{i\in[0,\ell):x_{i}\neq x^{(n)}_{i}\right\}\right|&\leq\ell_{n+1}+2\ell_{k}\cdot 5^{-n}+3\ell_{k}\cdot 5^{-n}+|u_{k}|+|E_{k}\cap[\ell_{k},\ell)|\\ &\leq\ell_{n+1}+\ell_{k}5^{-n+1}+|u_{k}|+|E_{k}\cap[\ell_{k},\ell)|.\end{split}$ (20)
Lastly, by the conditions (2) and (5) for the choice of $\ell_{k}$, it follows from (20) that
|$\begin{split}\frac{\left|\left\{i\in[0,\ell):x_{i}\neq x^{(n)}_{i}\right\}\right|}{\ell}&\leq\frac{\ell_{n+1}}{\ell}+\frac{\ell_{k}\cdot 5^{-n+1}}{\ell}+\frac{|u_{k}|}{\ell}+\frac{|E_{k}\cap[\ell_{k},\ell)|}{\ell}\\ &\leq\frac{\ell_{n+1}}{\ell}+5^{-n+1}+5^{-k}+5^{-k+1}.\end{split}$ (21)
Hence, $\bar{d}(x,x^{(n)})\leq 5^{-n+1}$ for all $n\in\mathbb{N}$, and $\lim_{n\to\infty}\bar{d}(x,x^{(n)})=0$. ∎
Corollary 6.3.
The proximal shift space $Z$ has the asymptotic average shadowing property.
6.2 A class of minimal shift spaces
In [6], the authors provided a way of constructing minimal shift spaces that are mixing and have the $\bar{d}$-shadowing property, acknowledging Piotr Oprocha for the ideas that inspired this construction. It is worth it mentioning that the obtained shift spaces have positive topological entropy (for definition and details we refer [6]). We will show that those shift spaces are $\bar{d}$-complete.
The parameters of the construction are an initial finite non-empty set of words ${\mathcal{B}}_{1}\subset\mathscr{A}^{*}$ and a sequence $\{t(n)\}_{n=1}^{\infty}\subset\mathbb{N}$ satisfying $t(n)\geq 2$ for all $n\in\mathbb{N}$.
Assume that we have defined the family of words ${\mathcal{B}}_{n}$ for some $n\in\mathbb{N}$. Write $k(n)$ for the cardinality of ${\mathcal{B}}_{n}$. Enumerate the elements of ${\mathcal{B}}_{n}$ as $\beta^{(n)}_{1},\ldots,\beta^{(n)}_{k(n)}$, and let $\tau(n)=\beta^{(n)}_{1}\ldots\beta^{(n)}_{k(n)}$ denote their concatenation. Let $s(n)$ (resp. $l(n)$) be the length of the shortest (resp. longest) word in ${\mathcal{B}}_{n}$. Words belonging to ${\mathcal{B}}_{n+1}$ are constructed as follows: first we concatenate $t(n)$ words arbitrarily chosen from ${\mathcal{B}}_{n}$, then we add the suffix $\tau(n)$, that is,
${\mathcal{B}}_{n+1}=\left\{b_{1}b_{2}\dots b_{t(n)}\tau(n):b_{i}\in{\mathcal{B}}_{n}\text{ for all }i\in[1,t(n)]\right\}.$
By construction, it holds
$\begin{split}s(n+1)&=t(n)s(n)+|\tau(n)|;\\ l(n+1)&=t(n)l(n)+|\tau(n)|.\end{split}$ (22)
For $n\in\mathbb{N}$, let $X_{n}$ be the coded shift space generated by the code ${\mathcal{B}}_{n}$. That is, $X_{n}$ consists of all infinite concatenations of words from ${\mathcal{B}}_{n}$ together with their shifts. The proofs of the facts below can be found in Section 4.2 of [6].
Fact 6.4.
Every word in ${\mathcal{B}}_{n}$ is a subword of every word in ${\mathcal{B}}_{m}$, for every $n<m$.
Fact 6.5.
The shift space $X_{n}$ is transitive and sofic for every $n\in\mathbb{N}$;
Fact 6.6.
For every $n\in\mathbb{N}$ it holds $X_{n+1}\subseteq X_{n}$, which implies $\bigcap_{n=1}^{\infty}X_{n}\neq\emptyset$.
Fact 6.7.
The shift space $X=\bigcap_{n=1}^{\infty}X_{n}$ is minimal.
From now on, we set ${\mathcal{B}}_{1}=\{0,11\}$. As noted in [6], from this choice of ${\mathcal{B}}_{1}$, a simple inductive argument shows that, for every $n\in\mathbb{N}$, the set of lengths of all words in ${\mathcal{B}}_{n}$ is an interval.
Proposition 6.8.
For every $m\in[s(n),l(n)]$, there is $u\in{\mathcal{B}}_{n}$ such that $|u|=m.$
Furthermore, we observe that the ratio between $s(n)$ and $l(n)$ is increasing. In particular, for all $n\in\mathbb{N}$ we have
$\begin{split}\dfrac{s(n)}{l(n)}\geq\dfrac{1}{2},\\ \dfrac{s(n)}{l(n)}<\dfrac{2}{3}.\end{split}$ (23)
The next lemma is a direct consequence of Proposition 6.8 and was extracted from the proof of Proposition 25 in [6]. For the sake of completeness we will replicate the argument.
Lemma 6.9.
Let $\alpha\in\mathbb{N}$ be such that $0<\alpha\leq l(n)$ and $b_{1},b_{2}\in{\mathcal{B}}_{n}$. Then one of the following holds:
  1. (1)
    There exist $c_{1},c_{2}\in{\mathcal{B}}_{n}$ such that $|c_{1}c_{2}|=|b_{1}b_{2}|+\alpha$;
  2. (2)
    There exist $c_{1},c_{2},c_{3}\in{\mathcal{B}}_{n}$ such that $|c_{1}c_{2}c_{3}|=|b_{1}b_{2}|+\alpha$.
Proof.
We first observe that $2s(n)<|b_{1}b_{2}|+\alpha\leq 3l(n)$. If $2s(n)<|b_{1}b_{2}|+\alpha\leq 2l(n)$, then, by Proposition 6.8, we can find $c_{1},c_{2}\in{\mathcal{B}}_{n}$ such that $|c_{1}c_{2}|=|b_{1}b_{2}|+\alpha$.
It follows from (23) that $3s(n)<2l(n)$. As a consequence, if $|b_{1}b_{2}|+\alpha>2l(n)$, then $3s(n)<|b_{1}b_{2}|+\alpha\leq 3l(n)$ and, by Proposition 6.8, we can find $c_{1},c_{2},c_{3}\in{\mathcal{B}}_{n}$ such that $|b_{1}b_{2}|+\alpha=|c_{1}c_{2}c_{3}|$. ∎
In [6], it is proved that under certain conditions on the sequence $\{t(n)\}_{n=1}^{\infty}\subset\mathbb{N}$ the obtained shift space $X$ possess various dynamical properties.
Theorem 6.10 (Theorem 27 in [6]).
There exist sequences of positive integers $\{t(n)\}_{n=1}^{\infty}$ such that $X$ is minimal and mixing, and has positive topological entropy and the $\bar{d}$-shadowing property.
The sequences $\{t(n)\}_{n=1}^{\infty}$ mentioned on Theorem 6.10 fulfill the following conditions for every $n\in\mathbb{N}$:
  1. (1)
    $t(n)>\frac{|\tau(n)|}{2l(n)-3s(n)}$;
  2. (2)
    $t(n)\geq\frac{l(n)}{l(n)-s(n)}$;
  3. (3)
    $t(n)\geq\frac{2s(n)+2l(n)+3|\tau(n)|}{l(n)}$;
  4. (4)
    $t(n)>\frac{3l(n)+|\tau(n)|}{s(n)2^{-n}}$.
From now on, we will assume that the sequence $\{t(n)\}_{n=1}^{\infty}$ fulfills the conditions above. We will show that these under these conditions the obtained shift space $X$ is $\bar{d}$-complete.
Lemma 6.11.
For every $x\in X$ and $n\in\mathbb{N}$ there exists a word $w\in{\mathcal{B}}_{n+1}$ such that
|$|\{i\in[0,|w|):x_{i}\neq w_{i}\}|\leq 3l(n)+|\tau(n)|.$
Proof.
Let $x\in X$. Since $x\in X_{n}$, there exists $\alpha\in[0,l(n))$ such that $x_{[\alpha,\infty)}$ can be written as a concatenation of words from ${\mathcal{B}}_{n}$, that is, there exists $\{b_{i}\}_{i=1}^{\infty}\subset{\mathcal{B}}_{n}$ such that
$x_{[\alpha,\infty)}=b_{1}b_{2}\cdots.$
By Lemma 6.9, one of the following holds:
  1. (1)
    there exist $c_{1},c_{2}\in{\mathcal{B}}_{n}$ such that $|x_{[0,\alpha)}b_{1}b_{2}|=|c_{1}c_{2}|$;
  2. (2)
    there exist $c_{1},c_{2},c_{3}\in{\mathcal{B}}_{n}$ such that $|x_{[0,\alpha)}b_{1}b_{2}|=|c_{1}c_{2}c_{3}|.$
Take $w\in{\mathcal{B}}_{n+1}$ as
$w=\begin{cases}c_{1}c_{2}b_{3}\ldots b_{t(n)}\tau(n),&\text{ if }(1)\text{ holds,}\\ c_{1}c_{2}c_{3}b_{3}\ldots b_{t(n)-1}\tau(n),&\text{ otherwise.}\end{cases}$
By construction, in the worst case scenario we would have
|$|\{i\in[0,|w|):x_{i}\neq w_{i}\}|\leq|uc_{1}c_{2}c_{3}|+|\tau(n)|\leq 3l(n)+|\tau(n)|.\qed$
We can improve Lemma 6.11 to ensure that the prefix of the obtained word $w$ can be chosen, while also providing more precise estimates for the number of distinct letters from $x_{[0,|w|)}$ and $w$ when analyzing subwords.
Lemma 6.12.
Let $x\in X$ and $\1n\in\mathbb{N}\backslash\{1\}$. For every $u\in{\mathcal{B}}_{n}$ there exists $w\in{\mathcal{B}}_{n+1}$ so that $u$ is prefix of $w$, and if $\ell\in[|u|,|w|)$, then
|$|\{i\in[|u|,\ell):x_{i}\neq w_{i}\}|\leq\ell\cdot 2^{-n+2}.$
Proof.
Fix $u\in{\mathcal{B}}_{n}$. Since $x_{[|u|,\infty)}\in X$, by Lemma 6.11, there exists $v^{(2)}\in{\mathcal{B}}_{n}$ such that
|$\left|\left\{i\in[0,|v^{(2)}|):x_{|u|+i}\neq v^{(2)}_{i}\right\}\right|\leq 3l(n-1)+|\tau(n-1)|.$
Similarly, for each $j\in[3,t(n)]$ we can find $v^{(j)}\in{\mathcal{B}}_{n}$ so that
|$\left|\left\{i\in[0,|v^{(j)}|):x_{|uv^{(2)}\ldots v^{(j-1)}|+i}\neq v^{(j)}_{i}\right\}\right|\leq 3l(n-1)+|\tau(n-1)|.$
Set $w=uv^{(2)}\ldots v^{(t(n))}\tau(n)\in{\mathcal{B}}_{n+1}$. For each $\ell\in[|u|,|w|-|\tau(n)|)$ we take the largest $j\in[2,t(n))$ so that $\ell\geq|uv^{(2)}\ldots v^{(t(j))}|$ and, by construction, we have
|$\left|\left\{i\in[|u|,\ell):x_{i}\neq w_{i}\right\}\right|\leq j\cdot\Bigl{(}3l\left(n-1\right)+|\tau(n-1)|\Bigr{)}.$ (24)
Note that condition (4) implies $3l(n-1)+|\tau(n-1)|\leq t(n-1)s(n-1)\cdot 2^{-n+1}$. Since $\ell\geq j\cdot s(n)$, it follows from (24) that
|$\begin{split}|\{i\in[|u|,\ell):x_{i}\neq w_{i}\}|&\leq j\cdot t(n-1)s(n-1)\cdot 2^{-n+1}\\ &\leq j\cdot s(n)\cdot 2^{-n+1}\\ &\leq\ell\cdot 2^{-n+1}.\end{split}$ (25)
For $\ell\in[|w|-|\tau|,|w|)$, note that condition (4) implies $|\tau(n)|<t(n)s(n)\cdot 2^{-n}$. Therefore, since $\ell\geq t(n)s(n)$, we conclude that
|$\begin{split}|\{i\in[|u|,\ell):x_{i}\neq w_{i}\}|&\leq t(n)s(n)\cdot 2^{-n+1}+|\tau(n)|\\ &\leq t(n)s(n)\cdot 2^{-n+1}+t(n)s(n)\cdot 2^{-n}\\ &\leq\ell\cdot 2^{-n+2}.\qed\end{split}$
Using a simple inductive argument we can extend Lemma 6.12 to the following result.
Lemma 6.13.
Let $x\in X$ and $n\in\mathbb{N}$. For every $u\in{\mathcal{B}}_{n}$ and $m>n$, there exists $w\in{\mathcal{B}}_{m}$ so that $u$ is prefix of $w$ and if $\ell\in[|u|,|w|)$, then
|$|\{i\in[|u|,\ell):x_{i}\neq w_{i}\}|\leq\ell\cdot 2^{-n+2}.$
Theorem 6.14.
The shift space $X$ is $\bar{d}$-complete.
Proof.
Let $\{x^{(n)}\}_{n=1}^{\infty}\subset X$ be a $\bar{d}$-Cauchy sequence. Up to pass to a subsequence we can assume that for every $n\in\mathbb{N}$ it holds
$\bar{d}\left(x^{(n)},x^{(n+1)}\right)<2^{-n-2}.$
We can construct a strictly increasing sequence $\{k_{n}\}_{n=1}^{\infty}\subset\mathbb{N}$ such that for every $n\in\mathbb{N}$ and $\ell\geq k_{n}$ it holds
|$\left|\left\{i\in[0,\ell):x^{(n)}_{i}\neq x^{(n+1)}_{i}\right\}\right|<\ell\cdot 2^{-n-2}.$ (26)
By the triangle inequality, it follows from (26) that for every $m>n$ and $\ell\geq k_{m-1}$ we have
|$\left|\left\{i\in[0,\ell):x_{i}^{(n)}\neq x_{i}^{(m)}\right\}\right|\leq\ell\cdot\left(2^{-n-2}+\ldots+2^{-m}\right)<\ell\cdot 2^{-n-1}.$ (27)
We will construct a strictly increasing sequence $\{m_{j}\}_{j=1}^{\infty}\subset\mathbb{N}$ and a sequence of words $\{w^{(j)}\}_{j=1}^{\infty}\subset\mathcal{L}(X)$ such that for all $j\in\mathbb{N}$ we have $w^{(j)}\in{\mathcal{B}}_{m_{j}}$, and $w^{(j)}$ as a prefix of $w^{(j+1)}$. We will denote $\alpha_{j}=|w^{(j)}|$. We then define $w\in\mathscr{A}^{\infty}$ to be the unique point such that for every $j\in\mathbb{N}$ we have $w_{[0,\alpha_{j})}=w^{(j)}$.
Let $m_{1}\geq 4$ be such that $s(m_{1})>k_{1}$. By Lemma 6.11, there exists $w^{(1)}\in B_{m_{1}}$ so that
|$\left|\left\{i\in[0,\alpha_{1}):x^{(1)}_{i}\neq w^{(1)}_{i}\right\}\right|\leq 3l(m_{1}-1)+\left|\tau(m_{1}-1)\right|.$
The construction follows by taking $m_{j+1}>m_{j}$ satisfying $s(m_{j+1})>k_{j+1}$. Note that $m_{1}\geq 4$ implies $m_{j}\geq j+3$ for all $j\in\mathbb{N}$. By Lemma 6.12, there exists $w^{(j+1)}\in{\mathcal{B}}_{m_{j+1}}$ with $w^{(j)}$ as a prefix, and so that for every $\ell\in[\alpha_{j},\alpha_{j+1})$ it holds
|$\left|\left\{i\in[\alpha_{j},\ell):x^{(j+1)}_{i}\neq w^{(j+1)}_{i}\right\}\right|\leq\ell\cdot 2^{-m_{j}+2}\leq\ell\cdot 2^{-j-1}.$ (28)
Note that, since $\{m_{j}\}_{j=1}^{\infty}\nearrow\infty$ and $w_{[0,\alpha_{j})}\in{\mathcal{B}}_{m_{j}}$, we have $w\in X$. It remains to show that $\lim_{n\to\infty}\bar{d}(x^{(n)},w)=0$. Fix $n\in\mathbb{N}$. For every $\ell\geq\alpha_{n+1}$ there exists $k=k(\ell)\in\mathbb{N}$ such that $\ell\in[\alpha_{n+k},\alpha_{n+k+1})$. By the triangle inequality, we have
|$\left|\left\{i\in[0,\ell):x^{(n)}_{i}\neq w_{i}\right\}\right|$ |$\leq\left|\left\{i\in[0,\alpha_{n}):x^{(n)}_{i}\neq w_{i}\right\}\right|$
$+\sum_{j=0}^{k-1}\left|\left\{i\in[\alpha_{n+j},\alpha_{n+j+1}):x^{(n)}_{i}\neq x^{(n+j+1)}_{i}\right\}\right|$
$+\sum_{j=0}^{k-1}\left|\left\{i\in[\alpha_{n+j},\alpha_{n+j+1}):x^{(n+j+1)}_{i}\neq w_{i}\right\}\right|$
+|$+\left|\left\{i\in[\alpha_{n+k},\ell):x^{(n)}_{i}\neq x^{(n+k+1)}_{i}\right\}\right|$
+|$+\left|\left\{i\in[\alpha_{n+k},\ell):x^{(n+k+1)}_{i}\neq w_{i}\right\}\right|.$
It is not hard to see that $\alpha_{n+j+1}\geq 2\alpha_{n+j}$ for all $j\in\mathbb{N}$. Moreover, by construction, we have $\alpha_{n+j}>k_{n+j}$ for every $j\in\mathbb{N}$. Thus, using (27) we obtain
$\sum_{j=0}^{k-1}\left|\left\{i\in[\alpha_{n+j},\alpha_{n+j+1}):x^{(n)}_{i}\neq x^{(n+j+1)}_{i}\right\}\right|\leq\sum_{j=1}^{k}\alpha_{n+j}\cdot 2^{-n-1}\leq\alpha_{n+k}\cdot 2^{-n}.$ (29)
It also follows from (27) that
|$\left|\left\{i\in[\alpha_{n+k},\ell):x^{(n)}_{i}\neq x^{(n+k+1)}_{i}\right\}\right|\leq\ell\cdot 2^{-n-1}.$ (30)
Using (28) we obtain
$\begin{split}\sum_{j=0}^{k-1}\left|\left\{i\in[\alpha_{n+j},\alpha_{n+j+1}):x^{(n+j+1)}_{i}\neq w_{i}\right\}\right|&\leq\sum_{j=1}^{k}\alpha_{n+j}\cdot 2^{-n-j-1}\\ &\leq\alpha_{n+k}\cdot 2^{-n}.\end{split}$ (31)
Hence, combining (28), (29), (30) and (31), we conclude that
|$\begin{split}\left|\left\{i\in[0,\ell):x^{(n)}_{i}\neq w_{i}\right\}\right|&\leq\alpha_{n}+\alpha_{n+k}\cdot 2^{-n}+\alpha_{n+k}\cdot 2^{-n}+\ell\cdot 2^{-n-1}+\ell\cdot 2^{-n-k-1}\\ &\leq\alpha_{n}+\ell\cdot 2^{-n+2}.\end{split}$
Therefore, we conclude that $\bar{d}\left(x^{(n)},w\right)\leq 2^{-n+2}$ for all $n\in\mathbb{N}$, and consequently, $\lim_{n\to\infty}\bar{d}\left(x^{(n)},w\right)=0$. ∎
Corollary 6.15.
The minimal shift space $X$ has the asymptotic average shadowing property.
7 Applications
The equivalence of the vague specification and the asymptotic average shadowing properties allows us to use results already known about one property when dealing with the other. Consequently, we can address questions raised about systems exhibiting these properties.
Kamae showed in [11] that the specification property together with expansivity are sufficient conditions for the vague specification property. However, it is already known that the specification property alone implies the asymptotic average shadowing property (see [14, 16, 17]). Therefore, we can remove the assumption of expansivity from Kamae’s result.
Corollary 7.1.
If a topological dynamical system has the specification property, then it has the vague specification property.
The rest of the section is devoted to answer questions from [7] and [14].
Question 7.2 ([7]).
What is the relation between the vague and weak specification properties?
Question 7.3 (Question 10.2 in [14]).
Is there a nontrivial minimal topological dynamical system with the almost specification property or the (asymptotic) average shadowing property?
Question 7.4 (Question 10.3 in [14]).
Does the (asymptotic) average shadowing property imply the almost specification property? Does the average shadowing property imply the asymptotic average shadowing property?
Question 7.5 (Question 10.6 in [14]).
Is the (asymptotic) average shadowing property or the almost specification property inherited by factors?
Answer to Question 7.2
To answer this question we rely on results from [14] and [17]. We decided to skip the definitions of the weak specification and the almost specification property, and only invoke the results that are relevant to our purpose. We refer [16] for the definitions and further information.
Theorem 7.6 (Theorem 18 in [17]).
If a topological dynamical system has the weak specification property, then it has the asymptotic average shadowing property.
Immediately from Theorem 4.7 we obtain the following corollary.
Corollary 7.7.
If a topological dynamical system has the weak specification property, then it has the vague specification property.
The converse is not true and the almost specification property is used to prove it. Due to [14], it is known that the asymptotic average shadowing property (and consequently the vague specification property) is a consequence of the almost specification property.
Theorem 7.8 (Theorem 3.5 in [14]).
If a topological dynamical system has the almost specification property, then it has the asymptotic average shadowing property.
Corollary 7.9.
If a topological dynamical system has the almost specification property, then it has the vague specification property.
It is known that the weak specification and the almost specification properties are independent of each other. Indeed, Theorem 17 in [18] shows that there are topological dynamical systems with almost specification but without weak specification. In Section 6 of [18] there is a construction of a topological dynamical system with weak specification, but without almost specification (we also refer to [14, 16, 18, 19] for more details).
Therefore, Corollary 7.9 implies the following result, which together with Corollary 7.7, answers Question 7.2.
Corollary 7.10.
There are topological dynamical systems with the vague specification property but without the weak specification property.
Answer to Question 7.3
In Section 6.2 we present the construction of a class of minimal shift spaces with the $\bar{d}$-shadowing property from [6] (in particular, these shift spaces have the average shadowing property). By Corollary 6.15, every shift space in this class also has the asymptotic average shadowing property, which shows that Question 7.3 has a positive answer.
Answer to Question 7.4
The relation between the asymptotic average shadowing property and the almost specification is already clarified in [16].
A step toward an answer for the second part of Question 7.4 is given by Theorem 4.5 for surjective topological dynamical systems and by Corollary 5.17 for surjective shift spaces. Although we could not provide a definitive answer to the question, the lack of examples with the average shadowing property but not Besicovitch complete suggests a potential positive answer.
Answer to the Question 7.5
Theorem 4.7 also provides a partial answer to this question. Indeed, we can use Proposition 2 in [11], which says that the vague specification property is inherited by factors, to confirm that the asymptotic average shadowing property is inherited by factors.
Corollary 7.11.
Every factor of a topological dynamical system with the asymptotic average shadowing property has the asymptotic average shadowing property.
Acknowledgements
Research of MEC was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885 and NCN Polonez Bis grant No. 2022/47/P/ST1/00854; AT was supported by NCN Polonez Bis grant no. 2021/43/P/ST1/02885. The authors wish to express their gratitude to Dominik Kwietniak for suggesting the problem and providing insightful discussions.